OpenStudy (anonymous):

Natural logarithms . ln radicandx-8 = 5

OpenStudy (anonymous):



OpenStudy (anonymous):

Use definition of log to get \[e^5=x-8\]Then solve for x: \[x=e^5-8\]

OpenStudy (anonymous):

\[x \approx140.4131591026\]

OpenStudy (anonymous):

This is very useful (what I have called the definition): \[y=\log_{b} x\]is equivalent to \[b^y=x\]